If 1.500g of Vanadium metal reacts with oxygen gas to give 2.679g of Vanadium oxide, what is the empirical formula of the product?
1
Expert's answer
2017-11-27T11:20:07-0500
Calculate moles of vanadium reacted:
n(V)=(m(V))/(A_r (V))=(1.500 g)/(50.942 g/mol)=0.02945 mol
Calculate mass of O in oxide:
m(O)=m(oxide)-m(V)=2.679 g-1.500 g=1.179 g
Calculate moles of O in oxide:
n(O)=(m(O))/(A_r (O))=(1.179 g)/(15.999 g/mol)=0.07369 mol
The ration of moles:
V:O=0.02945 mol:0.07369 mol=(0.02945 mol)/(0.02945 mol):(0.07369 mol)/(0.02945 mol)=1:2.5=2:5
The formula of oxide is V2O5.